OpenStudy (anonymous):

I have an SAT question. The ratio of salt to pepper in a recipe is 2:5. If the amount of salt is increased by 50 percent and the amount of pepper is decreased by 20 percent, what is the new ratio of salt and pepper? I know the answer is 3:4. But I don't understand how. My teacher told me that I had to multiply 1.5 to the amount of salt and multiply 0.8 to the amount of pepper. How did he get these numbers?

OpenStudy (anonymous):

cant answer test questions

OpenStudy (anonymous):

against CoC

hero (hero):

@gjo4596, When posting questions, try to resist the urge to post the answer. There's no incentive for curious minds seeking to discover the answer on their own if you do that.

OpenStudy (anonymous):

Its a pratice test. And okay Hero

hero (hero):

@pascuzzoa, I'm pretty sure it's an SAT Practice exam she's referring to.

OpenStudy (anonymous):

lets say that you have 2x amount of salt and 5x amount of pepper in the first case. Now then in the second case you will have (2x+ 50% of 2x)= 3x of salt and (5x -20% of 5x) = 4x of pepper finally... Now there is your new ratio which is????????

OpenStudy (anonymous):

3:4 Do you know how my teacher got the 1.5 and 0.8? I think it was supposed to be some sort of shortcut

OpenStudy (anonymous):

see.. any quantity when increased by 50% = 0.5 it becomes 1.5 times its original.. so he/she said 1.5 for the first one.. because 2x * 1.5 = 3x similarly, when any quantity is decreased by 20% = 0.2 it becomes 0.8 times its original.. so he/she said 0.8 for the second one.. because 5x * 0.8 = 4x

OpenStudy (anonymous):

Oh I see. Thank you very much!!!

hero (hero):

What's funny is, All you have to do is this: Ratio of Original amount: \[\frac{\text{Salt}}{\text{Pepper}} = \frac{2}{5}\] Ratio of New Amount: \[\frac{\text{Salt}}{\text{Pepper}} = \frac{2 + 1}{5 - 1} = \frac{3}{4}\] You simply add 1 to 2 since 1 is 50% of 2 and subtract 1 from 5 since 1 is 20% of 5 You get 3/4. That's all you have to do.

hero (hero):

@gjo4596